# AST-T5: Structure-Aware Pretraining for Code Generation and Understanding

### v3

@@ 2 Related Work
Language models (LMs) extended their use from NLP to code understanding and generation. Encoder-only models generally excel in code understanding when finetuned with classifiers (Feng et al., 2020), while decoder-only models are optimized for code generation through their autoregressive nature (Chen et al., 2021a; Fried et al., 2023; Nijkamp et al., 2023). However, these models can falter outside their primary domains of expertise or require increased resources for comparable outcomes. Our work focuses on encoder-decoder models, aiming to efficiently balance performance in both understanding and generation tasks without excessive computational demands.
@@ 3.2 AST-Aware Segmentation
    We construct an array cost, where cost[i] denotes the number of AST-structure breaks that would occur if partitioning happened right after token ii. This array is populated by traversing the AST and incrementing cost[l..r - 1] by 1 for each span $[l,r]$ associated with an AST subtree.
@@ 3.4 AST-Aware Subtree Corruption
Except for the strategy used to select masked tokens and the segmentation strategy described in Section 3.2 , our approach adheres to the workflow described in Section $3.$ Once subtrees are selected for masking and replaced with sentinel tokens, the encoder processes this modified input. Subsequently, the decoder is tasked with reconstructing the original tokens within the masked subtrees. A side-by-side comparison between our approach and the Vanilla Span Corruption in T5 is presented in Figure 1.
@@ 4 Experimental Setup
AST-T5 is pretrained on a subset of The Stack Dedup corpus (Kocetkov et al., 2022), a near-deduplicated version of The Stack—a 3.1TB collection of permissively licensed source code from GitHub cutoff at April 2022, spanning 30 programming languages. For our experiments, AST-T5’s training involves Python, Java, C, C++, C#, and Markdown subsets, comprising a 588GB dataset with 93M code and natural language files.
Each file is first parsed into its AST using the Tree-Sitter multi-language parser, and then tokenized with byte-level Byte-Pair Encoding (BPE) using a 64k BPE token vocabulary. Following AST-Aware Segmentation, these files are partitioned into chunks of 1,024 tokens. Our model is pretrained using the AST-Aware Subtree Corruption objective for 524 billion tokens (1,024 tokens per sequence, 1,024 sequences per batch, and 500k steps). For each training example, we apply AST-Aware Subtree Corruption of it is code, or apply Vanilla T5 Span Corruption of it is natural language. For code, the threshold, $\theta$, is uniformly sampled from 5 to 100. Pretraining uses PyTorch, Fairseq [^2] https://github.com/facebookresearch/fairseq and FlashAttention (Dao et al., 2022) and is conducted on 8 nodes, each with 8x NVIDIA A100 40GB GPUs. Further pretraining hyperparameters are detailed in Section A.3.
@@
Table 1: Overview of our evaluation benchmarks about test set size, task type, and evaluation metric for each task. “Generation” tasks involve mapping natural language to code, “Transpilation” tasks involve translating code from one programming language to another, and “Understanding” tasks involve classifying code into categorical labels. For MBPP, we follow Nijkamp et al. (2023) and evaluate our model on the entire “sanitized” subset without few-shot prompts. For evaluation metrics, “Pass@1” indicates code execution on unit-tests provided in the benchmark using a single generated code per example, with reported pass rates. “EM” (Exact Match) evaluates textual equivalence without execution by comparing two canonicalized code pieces. “Acc” means accuracy in classification tasks. We omit “BLEU scores” because high BLEU values (>> 50) can still correspond to unexecutable or significantly flawed code (Lu et al., 2021), which is not useful in real-world applications. We also discuss evaluation results using the CodeBLEU (Ren et al., 2020) metric in Section A.5.
@@
We finetune AST-T5 on the training datasets of all downstream tasks, adhering to the methodology by Raffel et al. (2020). For the HumanEval task, which lacks its own training dataset, we use CodeSearchNet (Husain et al., 2020), aligning with the approach of Wang et al. (2023). The prompt templates for finetuning are constructed using the PromptSource framework (Bach et al., 2022). The finetuning takes 50k steps, with the peak learning rate set at 10% of the pretraining learning rate. All other hyperparameters from pretraining are retained without further adjustments, and we train only one finetuned model. During inference, rank classification is employed for code understanding tasks and beam search for generative tasks, following Sanh et al. (2021). We evaluate our model on the test set using five prompt templates for each task and report the average performance.
@@ 5.2 Main Results
Table 3: Results of AST-T5 on downstream tasks compared with reported results of established language models. Evaluation metrics align with those in Table 1. Our focus is primarily on models with similar sizes as AST-T5, specifically the “Base” models (110M to 230M parameters), while comparisons against larger models are depicted in Figure 3. Some models are either encoder-only or decoder-only and are thus not suited for certain tasks. These results are labeled with “N/A” in this table because they are not available in the literature.
@@
Figure 3: Visualizations of AST-T5’s performance on HumanEval and MBPP compared to other models compared to models exceeding 230M parameters. Each point on each scatter plot represents a model. The x-axis shows the parameter count in log-scale, while the y-axis shows the Pass@1 rate on HumanEval or MBPP in log-scale. Model open-source status is color-coded: blue for open-source and red for proprietary.
Table 3 shows AST-T5’s performance on downstream tasks compared with previously published results of similarly sized models, specifically those within the “Base” scale (110M to 230M parameters). Figure 3(a) and Figure 3(b) extends this comparison, comparing AST-T5 with larger models using the HumanEval benchmark and the MBPP benchmark, respectively. These results show that:
@@ 6 Conclusion and Future Work
Broader Impact
@@ A.3 Pretraining Hyperparameters
A.4 Evaluation Results on Multi-Lingual Code Generation
@@
Table 5 presents a comparative analysis of our AST-T5 model on Python and Java subsets of the multi-lingual HumanEval and MBXP benchmarks (Athiwaratkun et al., 2023). This analysis includes models such as BLOOM (BigScience, 2021), OPT (Zhang et al., 2022), and various configurations of CodeGen (Nijkamp et al., 2023), as reported in Athiwaratkun et al. (2023). Our results show AST-T5’s superior performance across all benchmarks compared to the CodeGen-multi-350M. Furthermore, AST-T5, having 226M parameters, outperforms larger counterparts like BLOOM-7.1B and OPT-13B.
A.5 Evaluation Results in CodeBLEU

### v4

@@ 2 Related Work
Language models (LMs) extended their use from NLP to code understanding and generation. Encoder-only models generally excel in code understanding when finetuned with classifiers (Feng et al., 2020), while decoder-only models are optimized for code generation through their autoregressive nature (Chen et al., 2021a; Fried et al., 2023; Nijkamp et al., 2023b). However, these models can falter outside their primary domains of expertise or require increased resources for comparable outcomes. Our work focuses on encoder-decoder models, aiming to efficiently balance performance in both understanding and generation tasks without excessive computational demands.
@@ 3.2 AST-Aware Segmentation
    We construct an array cost, where cost[i] denotes the number of AST-structure breaks that would occur if partitioning happened right after token $i$. This array is populated by traversing the AST and incrementing cost[l..r - 1] by 1 for each span $[l,r]$ associated with an AST subtree.
@@ 3.4 AST-Aware Subtree Corruption
Except for the strategy used to select masked tokens and the segmentation strategy described in Section 3.2 , our approach adheres to the workflow described in Section 3.3. Once subtrees are selected for masking and replaced with sentinel tokens, the encoder processes this modified input. Subsequently, the decoder is tasked with reconstructing the original tokens within the masked subtrees. A side-by-side comparison between our approach and the Vanilla Span Corruption in T5 is presented in Figure 1.
@@ 4 Experimental Setup
AST-T5 is pretrained on a subset of The Stack Dedup corpus (Kocetkov et al., 2022), a near-deduplicated version of The Stack—a 3.1TB collection of permissively licensed source code from GitHub cutoff at April 2022, spanning 358 programming languages. For our experiments, AST-T5’s training involves Python, Java, C, C++, C#, Markdown, and reStructuredText subsets, comprising a 588GB dataset with 93M code and natural language files.
Each file is first parsed into its AST using the Tree-Sitter multi-language parser, and then tokenized with byte-level Byte-Pair Encoding (BPE) using a byte-level BPE token vocabulary. Following AST-Aware Segmentation, these files are partitioned into chunks of 1,024 tokens. Our model is pretrained using the AST-Aware Subtree Corruption objective for 524 billion tokens (1,024 tokens per sequence, 1,024 sequences per batch, and 500k steps). For each training example, we apply AST-Aware Subtree Corruption of it is code, or apply Vanilla T5 Span Corruption of it is natural language. For code, the threshold, $\theta$, is uniformly sampled from 5 to 100. For text, the length of each masked span is uniformly sampled from 1 to 10. Pretraining uses PyTorch, Fairseq [^2] https://github.com/facebookresearch/fairseq and FlashAttention (Dao et al., 2022) and is conducted on 8 nodes, each with 8x NVIDIA A100 40GB GPUs. Further pretraining hyperparameters are detailed in Section A.3.
@@
Table 1: Overview of our evaluation benchmarks about test set size, task type, and evaluation metric for each task. “Generation” tasks involve mapping natural language to code, “Transpilation” tasks involve translating code from one programming language to another, and “Understanding” tasks involve classifying code into categorical labels. For MBPP, we follow Nijkamp et al. (2023b) and evaluate our model on the entire “sanitized” subset without few-shot prompts. For evaluation metrics, “Pass@1” indicates code execution on unit-tests provided in the benchmark using a single generated code per example, with reported pass rates. “EM” (Exact Match) evaluates textual equivalence without execution by comparing two canonicalized code pieces. “Acc” means accuracy in classification tasks. We omit “BLEU scores” because high BLEU values (>> 50) can still correspond to unexecutable or significantly flawed code (Lu et al., 2021), which is not useful in real-world applications. We also discuss evaluation results using the CodeBLEU (Ren et al., 2020) metric in Section A.6.
@@
We finetune AST-T5 on the training datasets of all downstream tasks, adhering to the methodology by Raffel et al. (2020). For the HumanEval task, which lacks its own training dataset, we use CodeSearchNet (Husain et al., 2020), aligning with the approach of Wang et al. (2023). The prompt templates for finetuning are constructed using the PromptSource framework (Bach et al., 2022). The finetuning takes 50k steps, with the peak learning rate set at 10% of the pretraining learning rate. All other hyperparameters from pretraining are retained without further adjustments, and we train only one finetuned model. During inference, rank classification is employed for code understanding tasks and beam search is used for generative tasks, following Sanh et al. (2021). For CodeXGLUE, we evaluate our model on the test set using five prompt templates for each task and report the average performance; for HumanEval and MBPP, we evaluate the top-1 generated output from beam search.
@@ 5.2 Main Results
Table 3: Results of AST-T5 on downstream tasks compared with reported results of established language models. Evaluation metrics align with those in Table 1. Our focus is primarily on models with similar sizes as AST-T5, specifically the “Base” models (100M to 300M parameters), while comparisons against larger models are depicted in Figure 3. Some models are either encoder-only or decoder-only and are thus not suited for certain tasks. These results are labeled with “N/A” in this table because they are not available in the literature.
@@
Figure 3: Visualizations of AST-T5’s performance on HumanEval and MBPP compared to other models compared to models exceeding 300M parameters. Each point on each scatter plot represents a model. The x-axis shows the parameter count in log-scale, while the y-axis shows the Pass@1 rate on HumanEval or MBPP in log-scale. Model open-source status is color-coded: blue for open-source and red for proprietary.
Table 3 shows AST-T5’s performance on downstream tasks compared with previously published results of similarly sized models, specifically those within the “Base” scale (100M to 300M parameters). Figure 3(a) and Figure 3(b) extends this comparison, comparing AST-T5 with larger models using the HumanEval benchmark and the MBPP benchmark, respectively. Additional results on EvalPlus are shown in Section A.4. These results show that:
@@ 6 Conclusion and Future Work
Acknowledgements
We thank the reviewers and metareviewers at ICML for their constructive feedback and support. This work is supported in part by gift from Meta, the U.S. National Science Foundation through grants IIS-1955488, IIS-2027575, ARO W911NF2110339, ONR N00014-21-1-2724, and DOE awards DE-SC0016260, DE-SC0021982. The AI training platform and computation resources supporting this work were provided by High-Flyer AI Fundamental Research Co. Ltd, Hangzhou, China. We thank Sida Wang, Chris Cummins, Volker Seeker, and Hugh Leather for their valuable discussions.
Impact Statement
@@ A.3 Pretraining Hyperparameters
A.4 Evaluation Results on EvalPlus
We extend our evaluation to include EvalPlus (Liu et al., 2023), a more rigorous benchmark that enhances the original HumanEval and MBPP datasets with a substantial number of additional test cases. EvalPlus is designed to provide a more accurate evaluation of the correctness of programs produced by LLMs.
For our tests on HumanEval+ and MBPP+, we use the same hyperparameters used in our evaluations of HumanEval and MBPP. It is important to note that the hyperparameter configurations used in our study are not directly comparable to those used for the models listed on the EvalPlus leaderboard [^3] https://evalplus.github.io/leaderboard.html. Our results are compared against established models including GPT-Neo, GPT-J, InCoder, and CodeGen-2 (Nijkamp et al., 2023a).
Table 5: Performance of AST-T5 on HumanEval+ and MBPP+ benchmarks, compared with reported numbers of language models listed on the EvalPlus leaderboard. The evaluation metric used is Pass@1.
#Params HumanEval+ MBPP+
GPT-Neo 2.7B 6.7 7.9
GPT-J 6B 11.0 12.2
InCoder-1.3B 1.3B 11.0 12.2
InCoder-6.7B 6.7B 12.2 15.9
CodeGen2-1B 1B 9.1 11.0
CodeGen2-3B 3B 12.8 15.9
CodeGen2-7B 7B 17.7 18.3
CodeGen2-16B 16B 16.5 19.5
AST-T5 (Ours) 277M 12.8 19.3
As shown in Table 5, our 277M-parameter AST-T5 outperforms larger models like InCoder-6.7B and CodeGen2-1B, showing the effectiveness and parameter efficiency of AST-T5.
A.5 Evaluation Results on Multi-Lingual Code Generation
@@
Table 6 presents a comparative analysis of our AST-T5 model on Python and Java subsets of the multi-lingual HumanEval and MBXP benchmarks (Athiwaratkun et al., 2023). This analysis includes models such as BLOOM (BigScience, 2021), OPT (Zhang et al., 2022), and various configurations of CodeGen (Nijkamp et al., 2023b), as reported in Athiwaratkun et al. (2023). Our results show AST-T5’s superior performance across all benchmarks compared to the CodeGen-multi-350M. Furthermore, AST-T5, having 277M parameters, outperforms larger counterparts like BLOOM-7.1B and OPT-13B.
A.6 Evaluation Results in CodeBLEU